The longer leg of a right triangle exceeds the shorter leg by 1 inch, and the hypotenuse exceeds the longer leg by 7 inches. Find the length of the legs. Round to the nearest tenth of an inch. Thank you. Amber
Answer by mananth(16949)

let the shorter leg be x
the longer leg is x+1
The hypotenuse is (x+1)+7
x^2+(x+1)^2=(x+8)^2
x^2+x^2+2x+1=X^2 +16x+64
x^2-14x-63=0
X1= +14+sqrt(196+252) / 2
x1= 17.6 inches shorter leg
x2= +14-sqrt(448)/2
-3.58 Which is not possible
18.6 is the longer leg.